# Are foundation models for computer vision good conformal predictors?

### v2

@@ G Few-Shot Adaptation
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.7492 3.27 10.02 4.23 0.072 0.059 0.074
DINO-B 0.7743 2.58 10.73 3.03 0.072 0.058 0.073
DINOv2-S 0.7948 1.87 6.66 2.12 0.073 0.057 0.071
DINOv2-B 0.8393 1.36 7.50 1.67 0.076 0.057 0.065
DINOv2-L 0.8624 1.21 6.77 1.50 0.079 0.057 0.063
DINOv2-G 0.8666 1.18 4.08 1.44 0.079 0.058 0.063
VICReg (ResNet-50) 0.7207 4.24 14.22 5.73 0.070 0.058 0.071
VICReg (ResNet-50x2) 0.7570 3.08 13.49 3.89 0.070 0.056 0.070
VICReg (ResNet-200x2) 0.7804 2.50 12.05 2.98 0.071 0.056 0.071
CLIP 0.7201 3.03 9.50 3.73 0.069 0.058 0.070
MetaCLIP 0.7580 2.39 9.43 2.84 0.479 0.535 0.467
Table 4: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on ImageNet.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.8711 1.36 1.68 1.53 0.023 0.010 0.017
DINO-B 0.9165 1.14 1.41 1.30 0.024 0.009 0.013
DINOv2-S 0.9140 1.13 1.41 1.29 0.024 0.009 0.011
DINOv2-B 0.9511 1.01 1.22 1.14 0.025 0.008 0.009
DINOv2-L 0.9848 0.95 1.03 1.01 0.020 0.008 0.009
DINOv2-G 0.9932 0.95 0.99 0.98 0.020 0.008 0.008
VICReg (ResNet-50) 0.8506 1.49 1.84 1.69 0.021 0.011 0.018
VICReg (ResNet-50x2) 0.8749 1.32 1.65 1.50 0.022 0.010 0.014
VICReg (ResNet-200x2) 0.8481 1.47 1.85 1.66 0.022 0.011 0.017
CLIP 0.8823 1.33 1.71 1.51 0.021 0.009 0.013
MetaCLIP 0.8926 1.22 1.59 1.41 0.903 0.933 0.928
Table 5: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on CIFAR-10.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.6668 3.46 5.64 4.29 0.050 0.037 0.055
DINO-B 0.7379 2.19 4.07 2.50 0.050 0.034 0.050
DINOv2-S 0.7515 2.06 4.06 2.40 0.049 0.035 0.050
DINOv2-B 0.8116 1.39 3.40 1.76 0.054 0.034 0.042
DINOv2-L 0.8885 1.02 2.37 1.35 0.061 0.035 0.036
DINOv2-G 0.9235 0.95 1.93 1.21 0.067 0.035 0.036
VICReg (ResNet-50) 0.6503 3.83 6.15 4.83 0.050 0.034 0.056
VICReg (ResNet-50x2) 0.6902 3.06 5.30 3.78 0.054 0.035 0.056
VICReg (ResNet-200x2) 0.6461 3.95 6.23 4.91 0.053 0.037 0.058
CLIP 0.6667 3.52 6.28 4.18 0.050 0.037 0.055
MetaCLIP 0.7015 2.79 5.25 3.27 0.683 0.769 0.652
Table 6: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on CIFAR-100.
@@
Figure 8: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on ImageNet. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 9: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on ImageNet. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
@@
Figure 10: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on CIFAR-10. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 11: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on CIFAR-10. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
@@
Figure 12: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on CIFAR-100. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 13: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on CIFAR-100. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) LAC
(b) APS
(c) RAPS
Figure 15: Difference in conformal set size (i.e., efficiency) when applying temperature scaling on APS and CLIP ($T=1.1$).
ECE ($\downarrow$) AvgSize ($\downarrow$) MCCC ($\uparrow$)
(×$\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
DINO-S 5.59/3.02 3.27/3.26 10.02/11.78 4.23/3.99 0.433/0.427 0.477/0.480 0.412/0.415
DINO-B 3.94/2.12 2.58/2.55 10.73/13.14 3.03/3.03 0.416/0.398 0.521/0.517 0.391/0.409
VICReg (RN-50) 3.12/1.90 4.24/4.19 14.22/16.6 5.73/5.49 0.472/0.468 0.566/0.570 0.453/0.456
VICReg (RN-200x2) 2.21/1.90 2.50/2.49 12.05/14.61 2.98/3.01 0.440/0.430 0.567/0.566 0.462/0.471
CLIP 2.70/1.63 3.03/3.01 9.50/11.11 3.73/3.53 0.434/0.441 0.556/0.564 0.418/0.414
Table 7: Quantitative impact of calibrating vision foundation models. Average size and minimum class conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively) for several models, with ImageNet as benchmark dataset.
Set size ($\downarrow$) Coverage CovGap ($\downarrow$)

### v3

@@ Appendix F Few-Shot Adaptation
Table 4: Summary of the models used. Models used along with the number of parameters, training scheme, modalities used, and architecture type.
Num. of parameters Training scheme Modalities Architecture
DINO-S 21,665,664 SSL Vision ViT
DINO-B 85,798,656 SSL Vision ViT
DINOv2-S 22,056,576 SSL Vision ViT
DINOv2-B 86,580,480 SSL Vision ViT
DINOv2-L 304,368,640 SSL Vision ViT
DINOv2-G 1,136,480,768 SSL Vision ViT
VICReg (ResNet-50) 23,508,032 SSL Vision CNN
VICReg (ResNet-50x2) 93,907,072 SSL Vision CNN
VICReg (ResNet-200x2) 250,128,512 SSL Vision CNN
MetaCLIP 149,620,737 Contrastive Vision-Language ViT
Phi 3.5 303,507,456 SSL+Supervised Vision-Language ViT
LLaVa 303,507,456 SSL+Supervised Vision-Language ViT
CLIP (ViT-B) 87,456,000 Contrastive Vision-Language ViT
CLIP (ViT-L) 303,966,208 Contrastive Vision-Language ViT
CLIP (ViT-H) 632,076,800 Contrastive Vision-Language ViT
CLIP (ConvNeXt) 88,221,824 Contrastive Vision-Language CNN
CLIP (ConvNeXt-L) 199,770,816 Contrastive Vision-Language CNN
Table 5: Linear probing performance and conformal metrics for CIFAR-10. Linear probing F1 score, and corresponding set size and MCCC for LAC, APS, and RAPS ($\alpha=0.05$).
Set size ($\downarrow$) MCCC ($\uparrow$)
F1 ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.8711 1.36 $\pm$ 0.00 1.68 $\pm$ 0.00 1.53 $\pm$ 0.00 0.885 $\pm$ 0.001 0.930 $\pm$ 0.001 0.913 $\pm$ 0.001
DINO-B 0.9165 1.14 $\pm$ 0.00 1.41 $\pm$ 0.00 1.30 $\pm$ 0.00 0.896 $\pm$ 0.001 0.932 $\pm$ 0.001 0.921 $\pm$ 0.001
DINOv2-S 0.9140 1.13 $\pm$ 0.00 1.41 $\pm$ 0.00 1.29 $\pm$ 0.00 0.899 $\pm$ 0.001 0.932 $\pm$ 0.001 0.927 $\pm$ 0.001
DINOv2-B 0.9511 1.01 $\pm$ 0.00 1.22 $\pm$ 0.00 1.14 $\pm$ 0.00 0.899 $\pm$ 0.001 0.934 $\pm$ 0.001 0.933 $\pm$ 0.001
DINOv2-L 0.9848 0.95 $\pm$ 0.00 1.03 $\pm$ 0.00 1.01 $\pm$ 0.00 0.893 $\pm$ 0.001 0.933 $\pm$ 0.001 0.933 $\pm$ 0.001
DINOv2-G 0.9932 0.95 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00 0.910 $\pm$ 0.001 0.936 $\pm$ 0.001 0.934 $\pm$ 0.001
VICReg (ResNet-50) 0.8506 1.49 $\pm$ 0.00 1.84 $\pm$ 0.00 1.69 $\pm$ 0.00 0.902 $\pm$ 0.001 0.925 $\pm$ 0.001 0.905 $\pm$ 0.001
VICReg (ResNet-50x2) 0.8749 1.32 $\pm$ 0.00 1.65 $\pm$ 0.00 1.50 $\pm$ 0.00 0.910 $\pm$ 0.001 0.930 $\pm$ 0.001 0.923 $\pm$ 0.001
VICReg (ResNet-200x2) 0.8481 1.47 $\pm$ 0.00 1.85 $\pm$ 0.00 1.66 $\pm$ 0.00 0.907 $\pm$ 0.001 0.925 $\pm$ 0.001 0.912 $\pm$ 0.001
MetaCLIP 0.8926 1.22 $\pm$ 0.00 1.59 $\pm$ 0.00 1.41 $\pm$ 0.00 0.903 $\pm$ 0.001 0.933 $\pm$ 0.001 0.928 $\pm$ 0.001
Phi 3.5 0.8761 1.33 $\pm$ 0.00 1.74 $\pm$ 0.00 1.51 $\pm$ 0.00 0.906 $\pm$ 0.001 0.928 $\pm$ 0.001 0.921 $\pm$ 0.001
LLaVa 0.9159 1.14 $\pm$ 0.00 1.46 $\pm$ 0.00 1.31 $\pm$ 0.00 0.902 $\pm$ 0.001 0.935 $\pm$ 0.001 0.928 $\pm$ 0.001
CLIP (ViT-B) 0.8823 1.33 $\pm$ 0.00 1.71 $\pm$ 0.00 1.51 $\pm$ 0.00 0.909 $\pm$ 0.001 0.931 $\pm$ 0.001 0.921 $\pm$ 0.001
CLIP (ViT-L) 0.9642 0.98 $\pm$ 0.00 1.15 $\pm$ 0.00 1.09 $\pm$ 0.00 0.877 $\pm$ 0.001 0.934 $\pm$ 0.001 0.932 $\pm$ 0.001
CLIP (ViT-H) 0.9519 1.01 $\pm$ 0.00 1.26 $\pm$ 0.00 1.16 $\pm$ 0.00 0.882 $\pm$ 0.001 0.933 $\pm$ 0.001 0.932 $\pm$ 0.001
CLIP (ConvNeXt) 0.8817 1.31 $\pm$ 0.00 1.69 $\pm$ 0.00 1.49 $\pm$ 0.00 0.904 $\pm$ 0.001 0.933 $\pm$ 0.001 0.922 $\pm$ 0.001
CLIP (ConvNeXt-L) 0.9470 1.02 $\pm$ 0.00 1.27 $\pm$ 0.00 1.18 $\pm$ 0.00 0.876 $\pm$ 0.001 0.932 $\pm$ 0.001 0.932 $\pm$ 0.001
Table 6: Linear probing performance and conformal metrics for CIFAR-100. Linear probing F1 score, and corresponding set size and MCCC for LAC, APS, and RAPS ($\alpha=0.1$).
Set size ($\downarrow$) MCCC ($\uparrow$)
F1 ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.6668 3.46 $\pm$ 0.01 5.64 $\pm$ 0.01 4.29 $\pm$ 0.02 0.678 $\pm$ 0.003 0.753 $\pm$ 0.003 0.682 $\pm$ 0.003
DINO-B 0.7379 2.19 $\pm$ 0.01 4.07 $\pm$ 0.01 2.50 $\pm$ 0.01 0.667 $\pm$ 0.003 0.772 $\pm$ 0.003 0.678 $\pm$ 0.003
DINOv2-S 0.7515 2.06 $\pm$ 0.01 4.06 $\pm$ 0.01 2.40 $\pm$ 0.02 0.689 $\pm$ 0.004 0.761 $\pm$ 0.004 0.700 $\pm$ 0.004
DINOv2-B 0.8116 1.39 $\pm$ 0.01 3.40 $\pm$ 0.01 1.76 $\pm$ 0.00 0.671 $\pm$ 0.003 0.778 $\pm$ 0.003 0.739 $\pm$ 0.003
DINOv2-L 0.8885 1.02 $\pm$ 0.01 2.37 $\pm$ 0.01 1.35 $\pm$ 0.00 0.661 $\pm$ 0.003 0.779 $\pm$ 0.003 0.771 $\pm$ 0.002
DINOv2-G 0.9235 0.95 $\pm$ 0.00 1.93 $\pm$ 0.01 1.21 $\pm$ 0.00 0.583 $\pm$ 0.004 0.776 $\pm$ 0.003 0.762 $\pm$ 0.003
VICReg (ResNet-50) 0.6503 3.83 $\pm$ 0.01 6.15 $\pm$ 0.01 4.83 $\pm$ 0.02 0.696 $\pm$ 0.004 0.776 $\pm$ 0.002 0.702 $\pm$ 0.003
VICReg (ResNet-50x2) 0.6902 3.06 $\pm$ 0.01 5.30 $\pm$ 0.01 3.78 $\pm$ 0.01 0.670 $\pm$ 0.004 0.771 $\pm$ 0.002 0.658 $\pm$ 0.004
VICReg (ResNet-200x2) 0.6461 3.95 $\pm$ 0.01 6.23 $\pm$ 0.02 4.91 $\pm$ 0.01 0.713 $\pm$ 0.003 0.769 $\pm$ 0.002 0.687 $\pm$ 0.003
MetaCLIP 0.7015 2.79 $\pm$ 0.01 5.25 $\pm$ 0.01 3.27 $\pm$ 0.01 0.683 $\pm$ 0.004 0.769 $\pm$ 0.003 0.652 $\pm$ 0.003
Phi 3.5 0.6796 3.30 $\pm$ 0.01 6.58 $\pm$ 0.02 3.94 $\pm$ 0.01 0.650 $\pm$ 0.004 0.751 $\pm$ 0.003 0.654 $\pm$ 0.003
LLaVa 0.7201 2.56 $\pm$ 0.01 5.31 $\pm$ 0.02 2.97 $\pm$ 0.01 0.701 $\pm$ 0.004 0.762 $\pm$ 0.003 0.694 $\pm$ 0.003
CLIP (ViT-B) 0.6667 3.52 $\pm$ 0.01 6.28 $\pm$ 0.02 4.18 $\pm$ 0.01 0.705 $\pm$ 0.003 0.771 $\pm$ 0.002 0.691 $\pm$ 0.003
CLIP (ViT-L) 0.8315 1.28 $\pm$ 0.00 3.11 $\pm$ 0.01 1.67 $\pm$ 0.00 0.644 $\pm$ 0.003 0.774 $\pm$ 0.003 0.738 $\pm$ 0.003
CLIP (ViT-H) 0.7973 1.53 $\pm$ 0.00 3.79 $\pm$ 0.01 1.91 $\pm$ 0.00 0.653 $\pm$ 0.004 0.761 $\pm$ 0.003 0.709 $\pm$ 0.004
CLIP (ConvNeXt) 0.6880 3.21 $\pm$ 0.00 6.23 $\pm$ 0.02 3.99 $\pm$ 0.01 0.692 $\pm$ 0.003 0.777 $\pm$ 0.003 0.648 $\pm$ 0.004
CLIP (ConvNeXt-L) 0.7858 1.66 $\pm$ 0.00 3.73 $\pm$ 0.01 1.97 $\pm$ 0.00 0.680 $\pm$ 0.004 0.768 $\pm$ 0.002 0.714 $\pm$ 0.003
Table 7: Linear probing performance and conformal metrics for ImageNet. Linear probing F1 score, and corresponding set size and MCCC for LAC, APS, and RAPS ($\alpha=0.1$).
Set size ($\downarrow$) MCCC ($\uparrow$)
F1 ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.7492 3.27 $\pm$ 0.01 10.02 $\pm$ 0.02 4.23 $\pm$ 0.02 0.433 $\pm$ 0.005 0.477 $\pm$ 0.006 0.412 $\pm$ 0.005
DINO-B 0.7743 2.58 $\pm$ 0.00 10.73 $\pm$ 0.02 3.03 $\pm$ 0.00 0.416 $\pm$ 0.005 0.521 $\pm$ 0.005 0.391 $\pm$ 0.005
DINOv2-S 0.7948 1.87 $\pm$ 0.00 6.66 $\pm$ 0.01 2.12 $\pm$ 0.00 0.370 $\pm$ 0.006 0.520 $\pm$ 0.005 0.384 $\pm$ 0.006
DINOv2-B 0.8393 1.36 $\pm$ 0.00 7.50 $\pm$ 0.02 1.67 $\pm$ 0.00 0.414 $\pm$ 0.005 0.476 $\pm$ 0.005 0.489 $\pm$ 0.005
DINOv2-L 0.8624 1.21 $\pm$ 0.00 6.77 $\pm$ 0.02 1.50 $\pm$ 0.00 0.326 $\pm$ 0.004 0.502 $\pm$ 0.005 0.470 $\pm$ 0.004
DINOv2-G 0.8666 1.18 $\pm$ 0.00 4.08 $\pm$ 0.01 1.44 $\pm$ 0.00 0.244 $\pm$ 0.005 0.458 $\pm$ 0.006 0.350 $\pm$ 0.005
VICReg (ResNet-50) 0.7207 4.24 $\pm$ 0.01 14.22 $\pm$ 0.03 5.73 $\pm$ 0.01 0.472 $\pm$ 0.004 0.566 $\pm$ 0.004 0.453 $\pm$ 0.005
VICReg (ResNet-50x2) 0.7570 3.08 $\pm$ 0.01 13.49 $\pm$ 0.03 3.89 $\pm$ 0.01 0.442 $\pm$ 0.004 0.549 $\pm$ 0.005 0.445 $\pm$ 0.004
VICReg (ResNet-200x2) 0.7804 2.50 $\pm$ 0.00 12.05 $\pm$ 0.03 2.98 $\pm$ 0.00 0.440 $\pm$ 0.004 0.567 $\pm$ 0.004 0.462 $\pm$ 0.004
MetaCLIP 0.7580 2.39 $\pm$ 0.00 9.43 $\pm$ 0.02 2.84 $\pm$ 0.00 0.479 $\pm$ 0.005 0.535 $\pm$ 0.005 0.467 $\pm$ 0.004
Phi 3.5 0.7325 2.93 $\pm$ 0.00 10.89 $\pm$ 0.02 3.50 $\pm$ 0.01 0.449 $\pm$ 0.005 0.535 $\pm$ 0.005 0.464 $\pm$ 0.005
LLaVa 0.8500 1.27 $\pm$ 0.00 3.83 $\pm$ 0.01 1.57 $\pm$ 0.00 0.424 $\pm$ 0.004 0.537 $\pm$ 0.005 0.506 $\pm$ 0.005
CLIP (ViT-B) 0.7201 3.03 $\pm$ 0.00 9.50 $\pm$ 0.02 3.73 $\pm$ 0.00 0.434 $\pm$ 0.005 0.556 $\pm$ 0.004 0.418 $\pm$ 0.006
CLIP (ViT-L) 0.8438 1.34 $\pm$ 0.00 4.40 $\pm$ 0.01 1.66 $\pm$ 0.00 0.358 $\pm$ 0.006 0.528 $\pm$ 0.005 0.462 $\pm$ 0.006
CLIP (ViT-H) 0.8393 1.40 $\pm$ 0.00 7.85 $\pm$ 0.02 1.75 $\pm$ 0.00 0.391 $\pm$ 0.007 0.553 $\pm$ 0.004 0.457 $\pm$ 0.006
CLIP (ConvNeXt) 0.7787 2.01 $\pm$ 0.00 7.92 $\pm$ 0.02 2.28 $\pm$ 0.00 0.406 $\pm$ 0.007 0.532 $\pm$ 0.005 0.454 $\pm$ 0.006
CLIP (ConvNeXt-L) 0.8165 1.56 $\pm$ 0.00 5.31 $\pm$ 0.01 1.88 $\pm$ 0.00 0.452 $\pm$ 0.006 0.517 $\pm$ 0.006 0.456 $\pm$ 0.006
Table 8: Quantitative impact of calibrating vision foundation models with temperature scaling. Average size and minimum class-conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively), with ImageNet as benchmark dataset.
[…]
($\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
DINO-S 2.57 3.27/3.26 10.02/11.78 4.23/3.99 0.433/0.427 0.477/0.480 0.412/0.415
DINO-B 1.82 2.58/2.55 10.73/13.14 3.03/3.03 0.416/0.398 0.521/0.517 0.391/0.409
DINOv2-S 1.34 1.87/1.87 6.66/8.34 2.12/2.19 0.370/0.351 0.520/0.538 0.384/0.409
DINOv2-B 1.04 1.36/1.36 7.50/10.46 1.67/1.77 0.414/0.406 0.476/0.482 0.489/0.498
DINOv2-L 1.18 1.21/1.21 6.77/9.67 1.50/1.59 0.326/0.336 0.502/0.522 0.453/0.487
DINOv2-G 1.56 1.18/1.18 4.08/5.69 1.44/1.50 0.244/0.243 0.458/0.484 0.350/0.375
VICReg (RN-50) 1.22 4.24/4.19 14.22/16.60 5.73/5.49 0.472/0.468 0.566/0.570 0.453/0.456
VICReg (RN-50x2) 0.18 3.08/3.09 13.49/16.06 3.89/3.80 0.442/0.430 0.549/0.557 0.445/0.446
VICReg (RN-200x2) 0.31 2.50/2.49 12.05/14.61 2.98/3.01 0.440/0.430 0.567/0.566 0.462/0.471
MetaCLIP 0.64 2.39/2.40 9.43/11.31 2.84/2.84 0.479/0.477 0.535/0.541 0.467/0.469
Phi 3.5 -0.44 2.93/2.92 10.89/12.82 3.50/3.38 0.449/0.446 0.535/0.540 0.464/0.468
LLaVa 0.98 1.27/1.27 3.83/4.94 1.57/1.63 0.424/0.426 0.537/0.554 0.506/0.517
CLIP (ViT-B) 1.07 3.03/3.01 9.50/11.11 3.73/3.53 0.434/0.441 0.556/0.564 0.418/0.414
CLIP (ViT-L) 0.22 1.34/1.34 4.40/5.68 1.66/1.74 0.358/0.368 0.528/0.546 0.462/0.470
CLIP (ViT-H) 0.21 1.40/1.41 7.85/9.55 1.75/1.83 0.391/0.392 0.553/0.584 0.457/0.470
CLIP (ConvNeXt) -0.25 2.01/2.01 7.92/9.74 2.28/2.36 0.406/0.406 0.532/0.560 0.454/0.461
CLIP (ConvNeXt-L) -0.22 1.56/1.56 5.31/6.69 1.88/1.96 0.452/0.450 0.517/0.542 0.456/0.476
(a) LAC
(b) APS
(c) RAPS
Figure 9: Difference in conformal set size (i.e., efficiency) when applying temperature scaling ($T=1.1$) on APS and CLIP (ViT-B).
Table 9: Quantitative impact of calibrating vision foundation models with histogram binning. Average size and minimum class-conditional coverage are reported between uncalibrated/calibrated models, with ImageNet as benchmark dataset.
[…]
($\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
DINO-S 5.59 3.27/135.41 10.02/312.12 4.23/312.15 0.433/0.422 0.477/0.375 0.412/0.375
DINO-B 3.94 2.58/101.94 10.73/269.62 3.03/269.67 0.416/0.417 0.521/0.388 0.391/0.388
DINOv2-S 2.71 1.87/63.19 6.66/186.38 2.12/186.44 0.370/0.427 0.520/0.404 0.384/0.404
DINOv2-B 2.85 1.36/27.33 7.50/87.56 1.67/87.68 0.414/0.372 0.476/0.414 0.489/0.414
DINOv2-L 2.99 1.21/8.86 6.77/31.93 1.50/32.06 0.326/0.311 0.502/0.391 0.453/0.391
DINOv2-G 3.66 1.18/6.25 4.08/20.88 1.44/20.98 0.244/0.280 0.458/0.386 0.350/0.385
VICReg (RN-50) 3.14 4.24/161.07 14.22/374.82 5.73/374.84 0.472/0.456 0.566/0.362 0.453/0.362
VICReg (RN-50x2) 2.34 3.08/130.26 13.49/322.95 3.89/323.00 0.442/0.450 0.549/0.397 0.445/0.397
VICReg (RN-200x2) 2.21 2.50/101.10 12.05/268.30 2.98/268.37 0.440/0.444 0.567/0.384 0.462/0.384
MetaCLIP 2.63 2.39/110.80 9.43/286.46 2.84/286.51 0.479/0.446 0.535/0.396 0.467/0.396
Phi 3.5 2.29 2.93/140.27 10.89/329.86 3.50/329.87 0.449/0.456 0.535/0.384 0.464/0.385
LLaVa 2.47 1.27/13.72 3.83/47.75 1.57/47.87 0.424/0.380 0.537/0.401 0.506/0.402
CLIP (ViT-B) 2.70 3.03/148.29 9.50/337.35 3.73/337.35 0.434/0.458 0.556/0.374 0.418/0.374
CLIP (ViT-L) 1.86 1.34/24.29 4.40/74.14 1.66/74.21 0.358/0.402 0.528/0.418 0.462/0.418
CLIP (ViT-H) 2.00 1.40/30.29 7.85/99.07 1.75/99.19 0.391/0.405 0.553/0.427 0.457/0.427
CLIP (ConvNeXt) 1.86 2.01/84.08 7.92/231.47 2.28/231.52 0.406/0.449 0.532/0.407 0.454/0.407
CLIP (ConvNeXt-L) 1.66 1.56/42.62 5.31/139.04 1.88/139.14 0.452/0.435 0.517/0.413 0.456/0.413
@@
Figure 10: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the DINO and DINOv2 models on ImageNet. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet-50)
(b) VICReg (ResNet-50x2)
(c) VICReg (ResNet-200x2)
(d) MetaCLIP
(e) Phi 3.5
(f) LLaVa
Figure 11: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the VICReg, MetaCLIP, Phi 3.5 and LLaVa models on ImageNet. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
(a) CLIP (ViT-B)
(b) CLIP (ViT-L)
(c) CLIP (ViT-H)
(d) CLIP (ConvNeXt)
(e) CLIP (ConvNeXt-L)
Figure 12: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the CLIP models on ImageNet. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
@@
Figure 13: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the DINO and DINOv2 models on CIFAR-10. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet-50)
(b) VICReg (ResNet-50x2)
(c) VICReg (ResNet-200x2)
(d) MetaCLIP
(e) Phi 3.5
(f) LLaVa
Figure 14: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the VICReg, MetaCLIP, Phi 3.5 and LLaVa models on CIFAR-10. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
(a) CLIP (ViT-B)
(b) CLIP (ViT-L)
(c) CLIP (ViT-H)
(d) CLIP (ConvNeXt)
(e) CLIP (ConvNeXt-L)
Figure 15: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the CLIP models on CIFAR-10. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
@@
Figure 16: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the DINO and DINOv2 models on CIFAR-100. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet-50)
(b) VICReg (ResNet-50x2)
(c) VICReg (ResNet-200x2)
(d) MetaCLIP
(e) Phi 3.5
(f) LLaVa
Figure 17: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the VICReg, MetaCLIP, Phi 3.5 and LLaVa models on CIFAR-100. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
(a) CLIP (ViT-B)
(b) CLIP (ViT-L)
(c) CLIP (ViT-H)
(d) CLIP (ConvNeXt)
(e) CLIP (ConvNeXt-L)
Figure 18: Impact of the temperature $\bm{T}$ on the ECE, set size, MCCC, and CovGap, for the CLIP models on CIFAR-100. $T=1$ indicates the uncalibrated model performance. Green vertical line indicates the model performance for the optimal temperature.
(a) $\text{ViT}_{\text{MetaCLIP}}$ vs $\text{ViT}_{\text{ImageNet}}$
(b) $\text{ViT}_{\text{DINO-S}}$ vs $\text{ViT}_{\text{ImageNet}}$
Figure 19: Difference in set size for APS between (a) $\text{ViT}_{\text{MetaCLIP}}$ and $\text{ViT}_{\text{ImageNet}}$ and between (b) $\text{ViT}_{\text{DINO-S}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
(a) DINOv2-B, -A
(b) DINOv2-B, -R
(c) DINOv2-B, -Sketch
(d) VICReg (RN-50x2), -A
(e) VICReg (RN-50x2), -R
(f) VICReg (RN-50x2), -Sketch
(g) CLIP (ViT-B), -A
(h) CLIP (ViT-B), -R
(i) CLIP (ViT-B), -Sketch
Figure 21: Domain shift analysis. Distribution of class-conditional coverages for (a, b, c) DINOv2-B, (d, e, f) VICReg (RN 50x2), and (g, h, i) CLIP (ViT-B). The distribution shift datasets are (a, d, g) ImageNet-A, (b, e, h) ImageNet-R, and (c, f, i) and ImageNet-Sketch. APS (left) and RAPS (right).
Set size ($\downarrow$) Coverage ($\uparrow$) MCCC ($\uparrow$)